Solve for (0<=x<=360)
sin2x-cosx=-1+2sinx
ans:0,30,150,360

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Solve for (0<=x<=360)
sin2x-cosx=-1+2sinx
-------------
sin(2x)-cos(x)=-1+2sin(x)
2sin(x)cos(x) - cos(x) = -1 + 2sin(x)
cos*(2sin - 1) = -1 + 2sin = -(2sin - 1)
cos(x) = -1
x = 180 degs

Answer by ikleyn(52858)

.
Solve for (0 <= x <= 360)
sin2x-cosx=-1+2sinx
-------------

          Step by step


sin(2x) - cos(x) = -1 + 2sin(x)

2sin(x)cos(x) - cos(x) = -1 + 2sin(x)

cos(x)*(2sin(x) - 1) = -1 + 2sin(x)

cos(x)*(2sin(x) - 1) - (-1 + 2sin(x)) = 0

(cos(x)-1)*(2sin(x)-1) = 0


Case 1:  cos(x)-1 = 0    --->  cos(x) = 1  --->  x = 0.


Case 2:  2sin(x) -1 = 0  --->  2sin(x) = 1  --->  sin(x) = 1/2  --->  x = 30 degrees or x = 150 degrees.


ANSWER.  There are 3 (three) solutions: x= 0 degrees;  x = 30 degrees and x= 150 degrees.

Solved (correctly).

----------------------

For this problem, and for many other problems,  if the restriction for the argument
is for one period,  it is usually  (ALWAYS)  should be written in the form   0 <= x < 360
and  NOT  in the form   0 <= x <= 360,  because  0  and  360  represents
the same geometric angle  (although different numbers).


This small detail - how do you write this restriction - in correct form   0 <= x < 360
or in incorrect form  0 <= x <= 360,  makes it clear from the first glance
whether you know the subject and are familiar with standard mathematical notations - or not.